|Cloud Service - Office 365 (View update)||View History|
Last Updated: Friday, June 19th, 2015 at 3:31 pm
Office 365 running
|DNS Servers (View update)||View History|
Last Updated: Thursday, March 7th, 2013 at 10:46 am
DNS Servers Running
We can happily report that all DNS servers are currently running.
|DSL/FTTC Platform (View update)||View History|
Last Updated: Friday, November 6th, 2020 at 12:00 pm
Issue affecting our Broadband Service
6th November - 11:30 Our Broadband supplier is currently aware of a wide-spread packet loss issue following an outage instance that occurred last night. This is being thoroughly investigated for a resolution for as soon as possible. Further updates will follow as and when they become available and we apologise for any inconvenience caused. 6th November - 13:00 Latest update from our supplier is: "Due to the unplanned outage last night we are experiencing issues. As a result all traffic has been routed through an alternative BGP route and is subsequently causing congestion on this secondary route. we are currently in conversation with a supplier to establish whether or not we can bring this primary route back online. Despite this we are seeing that services are restored, but with issues." Further updates will be made available as soon as we have them. 6th November - 14:00 Latest update from our supplier is: "Emergency Work 14:00 – 16:00 6/11/2020 Due to the ongoing issues being observed on the DSL platform causing heavy packet loss, NOC engineers will soon be carrying out remedial work. All broadband users should be considered at risk for the duration of this work." We apologize for any inconvenience and further updates to be provided. 6th November - 14:45 Latest update from our supplier is: "This work has now been completed. We will continue to monitor the DSL platform. Any users still experiencing residual issues are advised to reboot the equipment prior to contacting the support desk." Once again we apologize for any inconvenience caused by the issue.
|Network Infrastructure (View update)||View History|
Last Updated: Tuesday, August 18th, 2020 at 10:38 am
Power Outage in Equinix Harbour Exchange
At around 4:24 this morning a fire alarm/power outage occurred at the Harbour Exchange/LD8 data centre in London. This power issue affected a number of floors in the DC and as of 8:40, equipment was being moved across to alternative power infrastructure. This issue affected one of our redundant backbone links as well as peering sessions with a number of other service providers. As of 9:26 our core network resilience was restored. Whilst our core network rerouted traffic over to remaining circuits and all our internet traffic was peering via other exchanges, you may have experienced some disruption whilst traffic was automatically rerouted and reduced performance due to additional load across all providers at the internet exchanges. Reports on LINX that other providers had their power restored during the morning meant that our BGP peering sessions were all recovered by 11:30. A number of our customers with leased line connections using Talk Talk infrastructure were affected due to the power outage with these services being restored by 15:15.
|Private Circuits + Leased Lines (View update)||View History|
Last Updated: Monday, June 25th, 2018 at 8:00 pm
Issue affecting Private circuits
June 25th 2018 20:00 Engineers have resolved all issues affecting the network June 25th 2018 10:25 We are currently investigating a problem on our network affecting some customers. We apologise for the inconvenience.
|Servers + Co-lo (View update)||View History|
Last Updated: Tuesday, September 26th, 2017 at 10:50 am
Loss of Connectivity to VM Platform
September 26th 2017 09:30 – We are aware of an issue with connectivity to the VM Platform. We are currently investigating and hope to restore service as soon as possible. We apologise for the inconvenience. UPDATE 10:05 - We traced the issue to the HA firewalls. The firewalls have been restarted and connectivity is now restored. If you continue to experience problems, please contact the support department (email@example.com).
|Veeam Cloud Connect (View update)||View History|
Last Updated: Wednesday, October 24th, 2018 at 4:20 pm
Issue Affecting Veeam Cloud Connect Backups
October 24 2018 13:00 Engineers have resolved all issues affecting the Veeam Cloud Connect service.